# Linying Lv

### v2

@@ 3.3.2 The Value of Analyst Information in the Cross-section of Analysts
Analysts who make bold forecasts incorporate their private information more completely and provide more relevant information to investors than those who herd in their forecasts (Clement and Tse, 2005). I examine how information value varies with analysts’ herding behavior.
To investigate the cross-section of analyst characteristics, I regress the stock-analyst-day information value on an indicator of forecast boldness, controlling for firm-fixed effects, analyst-fixed effects, and year-fixed effects. I also control for broker size, as Jacob et al. (1999) suggests that larger brokerage firms provide better research. The analyst-specific information value measure is defined as
@@
Table 10 presents the results. In columns (1)-(4) of Panel A, the coefficient estimates on $Bold$ are significantly positive, suggesting that analysts’ information is more strategically valuable when they provide bold forecasts. On average, information from bold reports is 29% more valuable than that from herding reports. This magnitude persists when examining the information value of report text alone. The insignificant coefficients for $Brokersize$ suggest that the value of analyst information is not related to brokerage firm size.
A subsequent research question is whether the relationship between analyst boldness and information value is driven by higher report informativeness or lower price impact. In Panel B, I regress the two decompositions of information value on the boldness indicator. As shown in columns (1)-(4), information from analysts who issue bold forecasts explains higher return volatility, supporting the argument of Clement and Tse (2005) that analysts provide bold forecasts when they possess relevant private information. In contrast, the coefficient estimates in columns (5)-(8), which examine the relationship between boldness and price impact, are statistically insignificant.
@@ 3.3.3 The Value of Analyst Information around Earnings Announcements
Previous evidence underlines the information content of analyst reports in the first week following earnings announcements. Accordingly, I evaluate the information value of analyst reports issued promptly following earnings announcements and beyond that.
Figure 5 illustrates the information value of analyst reports between earnings announcements on a weekly basis, from the first week after the earnings announcement up to the $13^{th}$ week. The information value of analyst reports is highest immediately following earnings announcements, particularly in the first week, peaking at a sample mean of \$0.84 million. This value sharply declines by the second week and stabilizes thereafter, with minor fluctuations. These patterns suggest that the market values analyst reports most highly immediately after earnings announcements, with diminishing importance in subsequent weeks.
To formally validate the observed pattern, I conduct a regression analysis of log information value on an indicator variable $Week$ which takes the value of one if the report is released within one week following earnings announcements and its interaction with earnings announcement day trading volume ($TradingVolume$). This model is motivated by prior literature suggesting that analysts’ interpretative role becomes more prominent with the complexity of financial accounting information (Chen et al., 2010). The trading volume serves as a proxy for investor disagreement, following Banerjee (2011) and Beckmann et al. (2024). The interaction term ($Week\times TradingVolume$) captures how post-earnings timing and investor disagreement jointly affect analyst information value. Stock and year-fixed effects are included to control for time-invariant stock characteristics and market-wide trends.
Table 11 presents regression results showing the influence of earnings announcements and trading volume on information value. Panel A indicates that the information value is significantly higher for reports released within one week of earnings announcements ($Week$), with positive and significant coefficients across all models. The interaction between $Week$ and $TradingVolume$ also shows a significant positive effect, suggesting that the information value increases with higher trading volumes around earnings announcements. The rationale is that investors may rely more on analyst opinions when there is greater disagreement in the interpretation of earnings information.
Panel B reveals significant increases in both log return variance and log price impact following earnings announcements, as evidenced by positive coefficients for the $Week$ variable. Notably, the increase in the information value of analyst reports following earnings announcements is primarily attributed to higher explained return volatility, rather than changes in price impact. These findings indicate that analyst reports provide more valuable insights in the first week after earnings announcements, especially during periods of high investor uncertainty. Investors value analysts’ ability to promptly interpret complex earnings information and make sense of the latest earnings numbers.
@@ 4 Conclusion
In this paper, I examine both quantitative and qualitative content of analyst reports using $R^{2}_{\mathrm{OOS}}$. The results show that report text contains more market-valued information content than quantifiable forecast revisions economically, especially right following earnings announcements. A Shapley value decomposition method reveals that analysts’ interpretation of realized income statements is the most valuable content within written reports. The text embedding representation extracts more comprehensive information from analyst reports beyond text sentiment.
I further translate the information content into dollar terms to provide a sense of economic magnitude. Using a measure of explained return volatility relative to price impact, I find that the annualized expected profit from early information acquisition of the analyst reports for an average S&P 100 constituent firm is larger than \$6.89 million. While these results are economically meaningful, readers should exercise caution in generalizing beyond the S&P 100 sample. Cross-sectional analysis reveals that the information value from analysts is higher for large stocks, bold forecasts, and reports released in the week following earnings announcements.
These findings have important implications for regulators, particularly concerning the fair distribution of information in financial markets. The significant potential profit from early access to analyst reports raises concerns about selective distribution. Such practices would violate FINRA Rule 2241, which aims to ensure the objectivity of research and prevent the selective distribution of research reports to preferred clients ahead of others. It is crucial to emphasize that this study presents a hypothetical scenario without assessing or endorsing actual market practices. Rather, by quantifying the potential profit of privileged access to analyst information, this research emphasizes the importance of regulations in ensuring fair disclosure.
@@ B. Additional Tables
Growth Strategies
Visa has been especially active on the acquisition front over the last several months. Combined with the Horizon assets and an emerging pipeline, there is enough in AMGN’s portfolio to offset potential headwinds and allow the company to grow through its patent cycle. Specific areas where more investment may be needed: Over the past decade, ROK management often claim that Process automation represents the growth opportunity for the company, but its sales in this market have barely grown in recent years, despite its much smaller sales base compared with established incumbents.
@@
Our DCF derives an intrinsic value of ¿\$100 for ABBV by discounting cash flows through 2024E, and assuming a -5% terminal growth & 7.6% WACC. We value MET based on a Sum-Of-The-Parts (SOTP) analysis based on our 2021 EPS estimate and using peer comps across each business segment. Our target price is based on a five-year discounted cash flow (DCF) valuation that employs a 5% discount rate and 20x terminal-year FCF multiple.
Investment Thesis
The results for the back half of the year will still be complex and confusing given the purchase accounting impacts and the full quarter of HSBC, but we do believe that there is a pay-off at the end of the road. As good as it gets: With its record multi-year backlog, Boeing’s revenue profile over the rest of decade is generally considered secure, and expectations for execution and cash already appear high. Clearly, our Ford Investment Thesis, which was based in large part on our belief that Ford would be able to offset headwinds (slowing cyclical tailwinds in North America, weakness in South America, weak growth in Europe, slowing growth in China, and regulatory cost headwinds), has been thrown into question.

### v3

@@ 3.3.2 The Value of Analyst Information in the Cross-section of Analysts
Prior literature suggests that analysts who make bold forecasts incorporate private information more comprehensively than those who herd (Clement and Tse, 2005). This section examines the relationship between information value and analysts’ herding behavior.
To analyze cross-sectional variation in analyst characteristics, I estimate regressions of stock-analyst-day information value on forecast boldness indicators, incorporating firm, analyst, and year fixed effects. Following Jacob et al. (1999), who suggests larger brokerage firms provide superior research, I include broker size as a control variable. The analyst-specific information value measure is defined as:
@@
Table 11 presents the regression results. Panel A, columns (1)-(4), reveals significantly positive coefficient estimates on $Bold$, indicating more strategically valuable information when analysts provide bold forecasts. Quantitatively, information from bold reports demonstrates 29% higher value compared to herding reports—a magnitude that persists when examining textual content alone. The insignificant coefficients for $Brokersize$ suggest that broker size is related to analyst information value.
To understand the underlying mechanisms, Panel B decomposes the relationship between analyst boldness and information value into informativeness and price impact components. As shown in columns (1)-(4), information from analysts who issue bold forecasts explains higher return volatility, supporting the argument of Clement and Tse (2005) that analysts issue bold forecasts when possessing relevant private information. Columns (5)-(8) reveal no statistically significant relationship between boldness and price impact, suggesting that the enhanced value of bold forecasts stems primarily from their information content rather than market microstructure effects.
@@ 3.3.3 The Value of Analyst Information around Earnings Announcements
Previous evidence underlines the information content of analyst reports in the first week following earnings announcements. Accordingly, this section evaluates the information value of analyst reports issued promptly following earnings announcements and beyond that.
Figure 5 examines different samples of analyst reports categorized by their timing relative to earnings announcements, from reports released in the first week after earnings announcements to those issued in the $13^{th}$ week. Reports issued in the first post-earnings week exhibit the highest information value, with a sample mean of \$0.84 million. Reports from subsequent weeks show a sharp decline in value. This pattern again suggests that markets place premium value on timely post-earnings analyst reports.
To formally test the observed pattern, I estimate regressions of log information value on an indicator variable $Week$, which identifies reports released within one week after earnings announcements. Following Banerjee (2011) and Beckmann et al. (2024), I include earnings announcement day trading volume ($TradingVolume$) as a proxy for investor disagreement. The model incorporates an interaction term ($Week\times TradingVolume$) to capture the joint effect of post-earnings timing and investor disagreement on analyst information value. The specification builds on Chen et al. (2010)’s finding that analysts’ interpretative role becomes more critical with increasing financial information complexity. The model includes stock and year fixed effects.
Table 12 presents regression results. Panel A documents significantly higher information value for reports issued within the first post-earnings week, as evidenced by positive and significant coefficients on $Week$ across all models. The positive and significant interactions between $Week$ and $TradingVolume$ indicate that information value increases with trading volume on earnings announcement days. The results are consistent with investors placing greater reliance on analyst interpretations during periods of heightened disagreement.
Panel B decomposes the earnings announcement effects into return variance and price impact, showing significant increases in both components. The coefficient magnitudes indicate that the timing effect is primarily driven by higher explained return volatility, rather than changes in price impact. These findings lead to conclude that investors value the most of analysts’ interpretation role during periods of heightened post-earnings uncertainty.
@@ 4 Conclusion
This paper examines both the quantitative and qualitative content of analyst reports. I find that report text contains more market-valued information content than quantifiable forecast revisions economically, especially shortly following earnings announcements. A Shapley value decomposition identifies analysts’ interpretation of income statements as the most valuable component of written reports.
I further translate the information content into economic terms. Using a measure of explained return volatility relative to price impact, I find that early acquisition of the analyst reports for an average S&P 100 constituent firm yields an annualized profit larger than \$6.89 million. Cross-sectional analysis reveals that the information value from analysts is higher for large stocks, bold forecasts, and reports released in the week following earnings announcements.
These findings have important implications for regulators, particularly concerning the fair distribution of information in financial markets. The substantial potential profits from early access to analyst reports raise concerns about selective distribution practices—a direct violation of FINRA Rule 2241, which mandates research objectivity and prohibits preferential distribution to certain clients. This study presents a hypothetical scenario for quantifying information value rather than an assessment of market practices. By documenting the magnitude of potential profits from privileged access, my research emphasizes the importance of regulations for fair disclosure.
@@ B. Additional Tables
Growth Strategies Visa has been especially active on the acquisition front over the last several months. Combined with the Horizon assets and an emerging pipeline, there is enough in AMGN’s portfolio to offset potential headwinds and allow the company to grow through its patent cycle. Specific areas where more investment may be needed: Over the past decade, ROK management often claim that Process automation represents the growth opportunity for the company, but its sales in this market have barely grown in recent years, despite its much smaller sales base compared with established incumbents.
@@
Our DCF derives an intrinsic value of \$100 for ABBV by discounting cash flows through 2024E, and assuming a -5% terminal growth & 7.6% WACC. We value MET based on a Sum-Of-The-Parts (SOTP) analysis based on our 2021 EPS estimate and using peer comps across each business segment. Our target price is based on a five-year discounted cash flow (DCF) valuation that employs a 5% discount rate and 20x terminal-year FCF multiple.
Investment Thesis The results for the back half of the year will still be complex and confusing given the purchase accounting impacts and the full quarter of HSBC, but we do believe that there is a pay-off at the end of the road. As good as it gets: With its record multi-year backlog, Boeing’s revenue profile over the rest of decade is generally considered secure, and expectations for execution and cash already appear high. Clearly, our Ford Investment Thesis, which was based in large part on our belief that Ford would be able to offset headwinds (slowing cyclical tailwinds in North America, weakness in South America, weak growth in Europe, slowing growth in China, and regulatory cost headwinds), has been thrown into question.